One side of a triangle is half the longest side. The third side is 13 feet less than the longest side. The perimeter is 72 feet. Find all three sides. (Enter your answers from smallest to largest.)
Answer by checkley77(12844)

x+x/2+x-13=72
x+.5x+x=72+13
2.5x=85
x=85/2.5
x=34 ans.
34+34/2+34-13=72
34+17+34-13=72
72=72